# Regge trajectories for the hidden bottom tetraquarks in the diquark picture

### v1

@@ Abstract
We apply the newly proposed Regge trajectory relation for the heavy-light diquarks alongside the Regge trajectory formula for the heavy-heavy systems to investigate the hidden bottom tetraquarks $(bq)(\bar{b}\bar{q}^{\prime})$ $(q,\,q^{\prime}=u,\,d,\,s)$. The Regge trajectory formulas for the hidden bottom tetraquarks are obtained. The masses of the $\lambda$-mode excited states and the $\rho$-mode excited states of the hidden bottom tetraquarks are estimated. Additionally, the $\lambda$-trajectories and the $\rho$-trajectories of the tetraquarks with hidden bottom are discussed. We find that the behavior of the $\rho$-trajectories is different from that of the $\lambda$-trajectories. The $\rho$-trajectories behave as $M{\sim}x_{\rho}^{1/2}$ $(x_{\rho}=n_{r},\,l)$ because the $\rho$-mode excitations are in diquark and antidiquark, with both the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$ being the heavy-light systems. Conversely, the $\lambda$-trajectories behave as $M{\sim}x_{\lambda}^{2/3}$ $(x_{\lambda}=N_{r},\,L)$ because the $\lambda$-mode excitations occure between diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$, with the the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$ forming a heavy-heavy system $(bq)(\bar{b}\bar{q}^{\prime})$. Moreover, not only the $\lambda$-trajectories but also the $\rho$-trajectories for the hidden bottom tetraquarks are concave downwards in the $(M^{2},\,x)$ plane.
Regge trajectory, tetraquark, mass
@@ I Introduction
As a type of exotic hadrons, tetraquarks will enhance the study of hadrons and will provide new probes for the understanding of QCD Jaffe:1976ig ; Jaffe:1976ih ; pdg:2024 ; Gross:2022hyw ; Bicudo:2022cqi ; Liu:2019zoy ; Jaffe:2004ph ; Esposito:2016noz . The hidden bottom tetraquarks, composed of two bottom quarks and two light quarks, have been studied through various methods. These approaches include the relativistic quark model based on the quasipotential approach in QCD Ebert:2008se ; Faustov:2021hjs , the non-relativistic potential model Tiwari:2022azj ; Lundhammar:2020xvw , the spinless Salpeter equation Ferretti:2020ewe ; Giannuzzi:2023vrx , the relativistic four-body Faddeev-Yakubovsky equation Hoffer:2024alv , the dynamical diquark model Giron:2020qpb , the homogeneous Lippmann-Schwinger integral equation Hadizadeh:2015cvx , the effective Hamiltonian in the adopted chromomagnetic interaction model Wu:2018xdi , the chiral quark model Liu:2022vyy , the sum rules Chen:2017dpy ; Wang:2019mxn , the relativistic four-quark equations in the framework of coupled-channel formalism Gerasyuta:2008cb , among others.
@@
The Regge trajectory is one of the effective approaches widely used in the study of hadron spectra Regge:1959mz ; Chew:1961ev ; Chew:1962eu ; Nambu:1978bd ; Collins:1971ff ; Inopin:1999nf ; Irving:1977ea . Ref. Selem:2006nd gives a compact and predictive unified picture of mesons, baryons, and tetraquarks with the loaded flux-tube model. Ref. Nielsen:2018uyn presents relations between the Regge trajectories of mesons, baryons, and tetraquarks based on the Holographic light front QCD. In Ref. Sonnenschein:2018fph , the masses of mesons, baryons, glueballs and tetraquarks are predicted by the Regge trajectories obtained from the Holography Inspired Stringy Hadron model. In Ref. MartinContreras:2020cyg , mesons, hybrid mesons and multiquark states are discussed by the nonlinear Regge trajectories in the model AdS/QCD with dilaton. In Refs. Chen:2023web ; Chen:2023djq , we discuss the universal descriptions of the heavy-heavy and heavy-light systems (mesons, baryons and tetraquarks) by using the Regge trajectories  [^1] A Regge trajectory of hadrons is assumed to be written as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$, where $M$ is mass of the bound state, $l$ is the the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, the plots in the $(M,\,x)$ plane, in the $(M,\,(x-c_{0})^{1/{\nu}})$ plane, in the $(M^{2},\,x)$ plane, in the $((M-m_{R})^{2},\,x)$ plane or in the $((M-m_{R})^{1/{\nu}},\,x)$ plane are all called the Chew-Frautschi plots. The Regge trajectories can be plotted in these different planes. . In the diquark picture, the hidden bottom tetraquarks are composed of one heavy-light diquark $(bq)$ and one heavy-light antidiquark $(\bar{b}\bar{q}^{\prime})$ ($q,q^{\prime}=u,d,s$). By applying the Regge trajectory relation for the heavy-light diquarks proposed in Ref. Chen:2023cws and the Regge trajectory formula for the heavy-heavy systems Chen:2023djq , we investigate the hidden bottom tetraquarks. The masses of the $\lambda$-mode excited and $\rho$-mode excited tetraquarks with the hidden bottom are estimated, and the $\lambda$-trajectories and $\rho$-trajectories are discussed.
@@
The paper is organized as follows: In Sec. II, the Regge trajectory relations are obtained from the spinless Salpeter equation. In Sec. III, we investigate the Regge trajectories for the hidden bottom tetraquarks. The conclusions are presented in Sec. IV.
II Regge trajectory relations for the hidden bottom tetraquarks
@@
In this section, with the help of the diquark Regge trajectories Chen:2023cws ; Feng:2023txx ; Chen:2023ngj , we present the Regge trajectory relations for the hidden bottom tetraquarks, which can discuss both the $\lambda$-trajectories and the $\rho$-trajectories.
@@ II.1 Preliminary
II.2 SSE
$\omega_{i}=\sqrt{m_{i}^{2}+{\bf p}^{2}}=\sqrt{m_{i}^{2}-\Delta}\;\;(i=1,2).$ (3)
II.3 Regge trajectory relations for the heavy-heavy systems
For the heavy-heavy systems, $m_{1},m_{2}{\gg}{{|\bf p|}}$, Eq. (2) reduces to
$+V_{d,t}\Psi_{d,t}({{\bf r}}),$
where
$\mu=m_{1}m_{2}/(m_{1}+m_{2}).$ (7)
By employing the Bohr-Sommerfeld quantization approach Brau:2000st and using Eqs. (4) and (6), the orbital and radial Regge trajectories can be obtained,
$M{\sim}\frac{3}{2}\left(\frac{\sigma^{\prime 2}}{\mu}\right)^{1/3}l^{2/3},\quad M{\sim}\left(\frac{3\pi}{2}\right)^{2/3}\left(\frac{\sigma^{\prime 2}}{2\mu}\right)^{1/3}{n_{r}}^{2/3},$ (8)
where
Consider the constant terms and the omitted terms and use Eq. (8), then we obtain Chen:2022flh ; Chen:2021kfw
$M=m_{R}+\beta_{x}(x+c_{0x})^{2/3}\,\,(x=l,\,n_{r}),$ (10)
with
$\beta_{x}=c_{fx}c_{x}c_{c},\quad m_{R}=m_{1}+m_{2}+C^{\prime},$ (11)
where
$c_{x}$ and $c_{c}$ are
$c_{c}=\left(\frac{\sigma^{\prime 2}}{\mu}\right)^{1/3},\quad c_{l}=\frac{3}{2},\quad c_{n_{r}}=\frac{\left(3\pi\right)^{2/3}}{2}.$ (13)
Both $c_{fl}$ and $c_{fn_{r}}$ are equal theoretically to one and are fitted in practice. In Eq. (10), $m_{1}$, $m_{2}$, $c_{x}$ and $\sigma$ are universal for the heavy-heavy systems. $c_{fx}$ and $c_{0x}$ are determined by fit.
II.4 Regge trajectory relations for the heavy-light systems
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , we obtain from Eq. (14)
$M{\sim}2\sqrt{\sigma^{\prime}}\sqrt{l},\;M{\sim}\sqrt{2{\pi}\sigma^{\prime}}\sqrt{n_{r}},$ (15)
where $\sigma^{\prime}$ is in Eq. (9). Using Eq. (15), the parameterized formula can be written as Chen:2022flh ; Chen:2021kfw
$M=m_{R}+\beta_{x}\sqrt{x+c_{0x}}\;(x=l,\,n_{r}).$ (16)
[In this subsection and the preceding subsection, the notations $m_{R}$, $\beta_{x}$, $c_{fx}$, $c_{x}$ and $c_{c}$ have the same form but different values. This should not cause confusion.] The parameter in Eq. (16) reads as
$\beta_{x}=c_{fx}c_{x}c_{c}$ (17)
with
$c_{c}=\sqrt{\sigma^{\prime}},\quad c_{l}=2,\quad c_{n_{r}}=\sqrt{2\pi}.$ (18)
Both $c_{fl}$ and $c_{fn_{r}}$ are theoretically equal to one but are fitted in practice. For the heavy-light systems, the common choice of $m_{R}$ is Selem:2006nd ; Chen:2021kfw ; Jakhad:2023mni ; Chen:2014nyo ; Chen:2017fcs ; Veseli:1996gy ; Jia:2018vwl
$m_{R}=m_{1}.$ (19)
$m_{R}$, $c_{x}$ and $\sigma$ are universal for the heavy-light systems. $c_{fx}$ and $c_{0x}$, which vary with different Regge trajectories, are determined by fitting the given Regge trajectory.
The usual Regge trajectory, Eq. (16) with (19), is obtained in the limit $m_{1}\to\infty$ and $m_{2}\to 0$. There are different ways of including the light constituent’s mass Selem:2006nd ; Nielsen:2018uyn ; Sonnenschein:2018fph ; MartinContreras:2020cyg ; Chen:2023cws ; Chen:2022flh ; Chen:2023ngj ; Chen:2014nyo ; Afonin:2014nya . Two modified formulas are proposed in Ref. Chen:2023cws , which can universally describe both the heavy-light mesons and the heavy-light diquarks. One is Eq. (16) with
$m_{R}=m_{1}+m_{2}+C^{\prime},$ (20)
where $C^{\prime}$ is in (12). Another reads
$M=m_{R}+\sqrt{\beta_{x}^{2}(x+c_{0x})+\kappa_{x}m^{3/2}_{2}(x+c_{0x})^{1/4}}$ (21)
if $m_{2}{\ll}M$, where
$m_{R}=m_{1}+C^{\prime},\quad\kappa_{x}=\frac{4}{3}\sqrt{{\pi}\beta_{x}},$ (22)
where $\beta_{x}$ is in (17). Eq. (16) with (20) is an extension of Afonin:2014nya
$M=m_{1}+m_{2}+\sqrt{a(n_{r}+{\alpha}l+b)},$ (23)
Table 1: The coefficients.
Heavy-heavy systems Heavy-light sytems
$\nu$ $2/3$ $1/2$
$c_{c}$ $\left({\sigma^{\prime 2}}/{\mu}\right)^{1/3}$ $\sqrt{\sigma^{\prime}}$
$c_{l}$ $3/2$ $2$
$c_{n_{r}}$ ${\left(3\pi\right)^{2/3}}/{2}$ $\sqrt{2\pi}$
In this work, Eq. (16) with (20) is employed to discuss the heavy-light systems. By summarising Eqs. (10) and (16) with (20), we have Chen:2022flh
$M=$ $m_{R}+\beta_{x}(x+c_{0x})^{\nu}\,\,(x=l,\,n_{r}),$
$m_{R}=$ $m_{1}+m_{2}+C^{\prime},\quad\beta_{x}=c_{fx}c_{x}c_{c},$ (24)
where ${\nu}$, $c_{x}$ and $c_{c}$ are listed in Table 1. $c_{fx}$ and $c_{0x}$ vary with different Regge trajectories. Eq. (II.4) can used to discuss the heavy-heavy and heavy-light systems: diquarks, mesons, baryons and tetraquarks Chen:2023web ; Chen:2023djq .
II.5 Regge trajectory relations for the hidden bottom tetraquarks
A hidden bottom tetraquark consists of one heavy-light diquark $(bq)$ and one heavy-light antidiquark $(\bar{b}\bar{q}^{\prime})$. Both $(bq)$ and $(\bar{b}\bar{q}^{\prime})$ are heavy because the bottom quark is heavy. Therefore, the hidden bottom tetraquarks are the heavy-heavy systems in case of the $\lambda$-mode excitations between diquark and antidiquark. Meanwhile, the Regge trajectories for hidden bottom tetraquarks exhibit the properties of the heavy-light systems in case of the $\rho$-mode excitations in diquark and antidiquark. According to discussions in the preceding subsections, using Eq. (II.4) and coefficients in Table 1, we obtain the Regge trajectory relation for the hidden bottom tetraquarks
@@ II.5 Regge trajectory relations for the hidden bottom tetraquarks
$\beta_{l_{1}}$ $=2\sqrt{\frac{\sigma}{2}}c_{fl_{1}},\quad\beta_{n_{r_{1}}}=\sqrt{\pi\sigma}c_{fn_{r_{1}}},$
$\beta_{l_{2}}$ $=2\sqrt{\frac{\sigma}{2}}c_{fl_{2}},\quad\beta_{n_{r_{2}}}=\sqrt{\pi\sigma}c_{fn_{r_{2}}}.$ (26)
There are three series of Regge trajectories for the hidden bottom tetraquarks, see Eq. (II.5).
III Regge trajectories for the hidden bottom tetraquarks
In this section, both the $\lambda-$trajectories and the $\rho-$trajectories for the hidden bottom tetraquarks are investigated.
III.1 Parameters
The quark masses, the string tension $\sigma$ and the parameter $C$ are from Ref. Faustov:2021hjs . The parameters for the heavy-light diquarks $(bu)$ and $(bs)$ are from Ref. Chen:2023cws and listed in Table 2. With these parameters determined, the $\rho-$modes and the diquark masses can be discussed, see Eqs. (II.5) and (II.5). To discuss the masses of the hidden bottom tetraquarks and the $\lambda-$modes excited states, the parameters $c_{fL}$, $c_{fN_{r}}$, $c_{0{L}}$ and $c_{0N_{r}}$ should be determined, see Eqs. (29) and (30). More discussions are in section C. Once all parameters are determined, the Regge trajectories can be discussed and can be used to estimate the masses of the hidden bottom tetraquarks.
@@ III.1 Parameters
$m_{u,d}=0.33\;{{\rm GeV}}$,   $m_{s}=0.50\;{{\rm GeV}}$,   $m_{b}=4.88\;{{\rm GeV}}$,
$\sigma=0.18\;{{\rm GeV}^{2}}$,  $C=-0.3\;{{\rm GeV}}$,
$(bu)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=0.988$,  $c_{fl_{1}}=c_{fl_{2}}=0.965$,
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.125$,
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.18$,
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.155$,
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.22$,
$(bs)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=0.953$,  $c_{fl_{1}}=c_{fl_{2}}=0.919$,
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.08$,
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.115$.
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.11$,
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.16$,
III.2 $\lambda-$ and $\rho-$trajectories
Table 3: The masses of the $\lambda$-mode excited states of the hidden bottom tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $n=u,\,d$. Eqs. (II.5), (29) and (30) are used.

### v4

@@ Abstract
We propose the Regge trajectory relations for the heavy tetraquarks $(Qq)(\bar{Q}\bar{q}^{\prime})$ $(Q=b,\,c;\,q,\,q^{\prime}=u,\,d,\,s)$ with hidden bottom and charm. By employing the new relations, both the $\lambda$-trajectories and the $\rho$-trajectories for the tetraquarks $(Qq)(\bar{Q}\bar{q}^{\prime})$ can be discussed. The masses of the $\lambda$-mode excited states and the $\rho$-mode excited states are estimated, and they agree with other theoretical predictions. We show that the behaviors of the $\rho$-trajectories are different from those of the $\lambda$-trajectories. The $\rho$-trajectories behave as $M{\sim}x_{\rho}^{1/2}$ $(x_{\rho}=n_{r},\,l)$ while the $\lambda$-trajectories behave as $M{\sim}x_{\lambda}^{2/3}$ $(x_{\lambda}=N_{r},\,L)$. Moreover, the Regge trajectory behaviors for other types of tetraquarks are investigated based on the spinless Salpeter equation. We show that both the $\lambda$-trajectories and the $\rho$-trajectories are concave downward in the $(M^{2},\,x)$ plane. The Regge trajectories for the tetraquarks containing the light diquark and/or the light antidiquark also are concave in the $(M^{2},\,x)$ plane when the masses of the light constituents are included and the confining potential is linear.
$\lambda$-trajectory, $\rho$-trajectory, tetraquark, mass
@@ I Introduction
As a type of exotic hadrons, tetraquarks will enhance the study of hadrons and will provide new probes for the understanding of QCD Jaffe:1976ig ; Jaffe:1976ih ; pdg:2024 ; Gross:2022hyw ; Bicudo:2022cqi ; Liu:2019zoy ; Jaffe:2004ph ; Esposito:2016noz . The heavy tetraquarks $(Qq)(\bar{Q}\bar{q}^{\prime})$ $(Q=b,\,c;\,q,q^{\prime}=u,\,d,\,s)$ with hidden bottom and charm have been studied through various methods. These approaches include the relativistic quark model based on the quasipotential approach in QCD Ebert:2008se ; Faustov:2021hjs ; Ebert:2008kb , the non-relativistic potential model Tiwari:2022azj ; Lundhammar:2020xvw , the spinless Salpeter equation Ferretti:2020ewe ; Giannuzzi:2023vrx , the relativistic four-body Faddeev-Yakubovsky equation Hoffer:2024alv , the dynamical diquark model Giron:2020qpb ; Lebed:2024zrp ; Giron:2020fvd ; Giron:2021sla , the homogeneous Lippmann-Schwinger integral equation Hadizadeh:2015cvx , the effective Hamiltonian in the adopted chromomagnetic interaction model Wu:2018xdi , the chiral quark model Liu:2022vyy , the sum rules Chen:2017dpy ; Wang:2019mxn , the relativistic four-quark equations in the framework of coupled-channel formalism Gerasyuta:2008cb , the constituent quark model Maiani:2004vq , among others.
@@
The Regge trajectory [^1] A Regge trajectory of hadrons is assumed to be written as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$, where $M$ is mass of the bound state, $l$ is the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, the plots in the $(M,\,x)$ plane Chen:2022flh ; Chen:2023cws , in the $(M,\,(x-c_{0})^{\nu})$ plane Burns:2010qq , in the $(M^{2},\,x)$ plane Chen:2018nnr ; Chen:2021kfw , in the $((M-m_{R})^{2},\,x)$ plane Chen:2023djq ; Chen:2023web or in the $((M-m_{R})^{1/{\nu}},\,x)$ plane [see Fig. 6 in this work] are all called the Chew-Frautschi plots. The Regge trajectories can be plotted in these different planes. is one of the effective approaches widely used in the study of hadron spectra Regge:1959mz ; Chew:1961ev ; Chew:1962eu ; Nambu:1978bd ; Collins:1971ff ; Inopin:1999nf ; Irving:1977ea . Ref. Selem:2006nd gives a compact and predictive unified picture of mesons, baryons, and tetraquarks with the loaded flux-tube model. Ref. Nielsen:2018uyn presents relations between the Regge trajectories of mesons, baryons, and tetraquarks based on the Holographic light front QCD. In Ref. Sonnenschein:2018fph , the masses of mesons, baryons, glueballs, and tetraquarks are predicted by the Regge trajectories obtained from the holography inspired stringy hadron model. In Ref. MartinContreras:2020cyg , mesons, hybrid mesons and multiquark states are discussed by the nonlinear Regge trajectories in the model AdS/QCD with dilaton. In Refs. Chen:2023web ; Chen:2023djq , we discuss the universal descriptions of the heavy-heavy and heavy-light systems by using the Regge trajectories. In preceding works, the Regge trajectories for the tetraquarks are the $\lambda$-trajectories. To our knowledge, no studies address the $\rho$-trajectories for tetraquarks although there are many studies on the $\rho$-excited states of tetraquarks. In the diquark picture, the hidden bottom tetraquarks are composed of one heavy-light diquark $(Qq)$ and one heavy-light antidiquark $(\bar{Q}\bar{q}^{\prime})$. By applying the Regge trajectory relation for the heavy-light diquarks Chen:2023cws and the Regge trajectory formula for the heavy-heavy systems Chen:2023djq , we present the Regge trajectory relations for the hidden bottom and charm tetraquarks. The masses of the $\lambda$-mode excited and $\rho$-mode excited tetraquarks with hidden bottom and charm are estimated. Both the $\lambda$-trajectories and $\rho$-trajectories are investigated.
@@
The paper is organized as follows: In Sec. II, the Regge trajectories for the hidden bottom and charm tetraquarks are discussed. The conclusions are presented in Sec. III. The discussions on the Regge trajectory behaviors for various tetraquarks are in the appendix C.
II Regge trajectories for the tetraquarks with hidden bottom and charm
@@
In this section, with the help of the diquark Regge trajectories Chen:2023cws ; Feng:2023txx ; Chen:2023ngj , we investigate the Regge trajectory behaviors for various tetraquarks. The Regge trajectory relations are presented, which can be employed to discuss both the $\lambda$-trajectories and the $\rho$-trajectories.
@@ II.1 Preliminary
II.2 Regge trajectory relations for the hidden bottom and charm tetraquarks
The tetraquarks with hidden bottom and charm consists of one heavy-light diquark $(Qq)$ and one heavy-light antidiquark $(\bar{Q}\bar{q}^{\prime})$. Both $(Qq)$ and $(\bar{Q}\bar{q}^{\prime})$ are heavy because the bottom/charm quark are heavy. Therefore, the tetraquarks $(Qq)(\bar{Q}\bar{q}^{\prime})$ are the heavy-heavy systems in case of the $\lambda$-mode excitations. Meanwhile, the Regge trajectories for $(Qq)(\bar{Q}\bar{q}^{\prime})$ exhibit the properties of the heavy-light systems in case of the $\rho$-mode excitations. According to Eqs. (C.3) and (C.3), we list the relations for the tetraquarks with hidden bottom and charm
@@ II.2 Regge trajectory relations for the hidden bottom and charm tetraquarks
$\beta_{l_{1}}$ $=\sqrt{2\sigma}c_{fl_{1}},\quad\beta_{n_{r_{1}}}=\sqrt{\pi\sigma}c_{fn_{r_{1}}},$
$\beta_{l_{2}}$ $=\sqrt{2\sigma}c_{fl_{2}},\quad\beta_{n_{r_{2}}}=\sqrt{\pi\sigma}c_{fn_{r_{2}}}.$ (3)
$M$, $M_{\rho_{1}}$, $M_{\rho_{2}}$, $m_{Q}$, and $m_{q}$ are the tetraquark mass, the diquark mass, the antidiquark mass, the heavy quark mass, and the light quark mass, respectively. $\sigma$ is the string tension. $C$ is a fundamental parameter. $c_{fx}$ are theoretically equal to one but are fitted in practice. $c_{0x}$ vary with Regge trajectories.
II.3 Parameters
The quark masses, the string tension $\sigma$ and the parameter $C$ are from Ref. Faustov:2021hjs . The parameters for the heavy-light diquarks $(bu)$, $(bs)$, $(cu)$, and $(cs)$ are from Ref. Chen:2023cws and listed in Table 1. With these parameters determined, the $\rho$-modes and the diquark masses can be discussed, see Eqs. (II.2) and (II.2). To discuss the masses of the hidden bottom and charm tetraquarks and the $\lambda$-modes excited states, the parameters $c_{fL}$, $c_{fN_{r}}$, $c_{0{L}}$ and $c_{0N_{r}}$ should be determined, see Eqs. (42) and (43). More discussions are in section D. Once all parameters are determined, the Regge trajectories can be discussed and can be used to estimate the masses of the tetraquarks with hidden bottom and charm.
@@ II.3 Parameters
$m_{u,d}=0.33\;{{\rm GeV}}$,   $m_{s}=0.50\;{{\rm GeV}}$,   $m_{b}=4.88\;{{\rm GeV}}$
$m_{c}=1.55\;{{\rm GeV}}$,   $\sigma=0.18\;{{\rm GeV}^{2}}$,  $C=-0.3\;{{\rm GeV}}$
$(bu)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=0.988$,  $c_{fl_{1}}=c_{fl_{2}}=0.965$
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.125$
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.18$
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.155$
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.22$
$(bs)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=0.953$,  $c_{fl_{1}}=c_{fl_{2}}=0.919$
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.08$
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.115$
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.11$
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.16$
$(cu)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=1.000$,  $c_{fl_{1}}=c_{fl_{2}}=1.038$
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.065$
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.095$
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.17$
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.19$
$(cs)$ $c_{fn_{r_{1}}}=c_{fn_{r_{2}}}=1.016$,  $c_{fl_{1}}=c_{fl_{2}}=1.015$
$c_{0n_{r_{1}}}(1^{1}s_{0})=c_{0n_{r_{2}}}(1^{1}s_{0})=0.03$
$c_{0n_{l_{1}}}(1^{1}s_{0})=c_{0n_{l_{2}}}(1^{1}s_{0})=0.055$
$c_{0n_{r_{1}}}(1^{3}s_{1})=c_{0n_{r_{2}}}(1^{3}s_{1})=0.095$
$c_{0n_{l_{1}}}(1^{3}s_{1})=c_{0n_{l_{2}}}(1^{3}s_{1})=0.135$
II.4 $\lambda$- and $\rho$-trajectories for the hidden bottom tetraquarks
Table 2: The masses of the $\lambda$-mode excited states of the hidden bottom tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $n=u,\,d$. Eqs. (II.2), (42) and (43) are used.